# MaskGCT: Zero-Shot Text-to-Speech with Masked Generative Codec Transformer

### v1

@@ 3.3 Speech Semantic Representation Codec
Discrete speech representations can be divided into semantic tokens and acoustic tokens. Generally, semantic tokens are obtained by discretizing features from speech self-supervised learning (SSL). Previous two-stage, large-scale TTS systems (Betker, 2023; Borsos et al., 2023; Kharitonov et al., 2023) typically first use text to predict semantic tokens, and then employ another model to predict acoustic tokens or features. This is because semantic tokens have a stronger correlation with text or phonemes, which makes predicting them more straightforward than directly predicting acoustic tokens. Commonly, previous works have used k-means to discretize semantic features to obtain semantic tokens; however, this method can lead to a loss of information. This loss may complicate the accurate reconstruction of high-quality speech or the precise prediction of acoustic tokens, especially for tonally rich languages. For example, our early experiments demonstrate the challenges of accurately predicting acoustic tokens to achieve proper prosody for Chinese using semantic tokens obtained via k-means. Therefore, we need to discretize semantic representation features while minimizing information loss. Inspired by Huang et al. (2023), we train an encoder-decoder model to learn a vector quantization codebook that reconstructs speech semantic representations from a speech SSL model (e.g., W2v-BERT (Chung et al., 2021)). For a speech semantic representation sequence $\mathbf{S}\in\mathbb{R}^{T\times d}$, the vector quantizer quantizes the output of the encoder $\mathcal{E}(\mathbf{S})$ to $\mathbf{E}$, and the decoder reconstructs $\mathbf{E}$ back to $\hat{\mathbf{S}}$. We optimize the encoder and the decoder using a reconstruction loss between $\mathbf{S}$ and $\hat{\mathbf{S}}$, employ codebook loss to optimize the codebook and use commitment loss to optimize the encoder with the straight-through method (Van Den Oord et al., 2017). The total loss for training the semantic representation codec can be written as:
@@
In detail, we utilize the hidden states from the 17th layer of W2v-BERT 2.0 (Chung et al., 2021) as the semantic features for our speech encoder. The encoder and decoder are composed of multiple ConvNext (Liu et al., 2022a) blocks. Following the methods of improved VQGAN (Yu et al., 2021) and DAC (Kumar et al., 2024), we use factorized codes to project the output of the encoder into a low-dimensional latent variable space. The codebook contains 8,192 entries, each of dimension 8. Further details about the model architecture are provided in the Appendix B.
3.4 Speech Acoustic Codec
Speech acoustic codec is trained to quantize speech waveform to multi-layer discrete tokens while aiming to preserve all the information of the speech as soon as possible. We follow the residual vector quantization (RVQ) method to compress the 24K sampling rate speech waveform into discrete tokens of 12 layers. The codebook size of each layer is 1,024 and the codebook dimension is 8. The model architectures, discriminators, and training losses follow DAC (Kumar et al., 2024), except that we use the Vocos (Siuzdak, 2023) architecture as the decoder for more efficient training and inference. Figure 5 shows the comparison between the semantic codec and acoustic codec.
3.5 Text-to-Semantic Model
@@
Based on the previous discussion, we employ a non-autoregressive masked generative transformer to train a text-to-semantic (T2S) model, instead of using an autoregressive model or any text-to-speech alignment information. During training, we randomly extract a portion of the prefix of the semantic token sequence as the prompt, denoted as $\mathbf{S}^{p}$. We then concatenate the text token sequence $\mathbf{P}$ with $\mathbf{S}^{p}$ to form the condition. We simply add $(\mathbf{P},\mathbf{S}^{p})$ as the prefix sequence to the input masked semantic token sequence $\mathbf{S}_{t}$ to leverage the in-context learning ability of language models. We use a Llama-style (Touvron et al., 2023) transformer as the backbone of our model, incorporating gated linear units with GELU (Hendrycks and Gimpel, 2016) activation, rotation position encoding (Su et al., 2024), etc., but replacing causal attention with bidirectional attention. We also use adaptive RMSNorm (Zhang and Sennrich, 2019), which accepts the time step $t$ as the condition.
@@
During inference, we generate the target semantic token sequence of any specified length conditioned on the text and the prompt semantic token sequence. In this paper, we also train a flow matching (Lipman et al., 2022) based duration prediction model to predict the total duration conditioned on the text and prompt speech duration, leveraging in-context learning. More details can be found in Appendix C.
3.6 Semantic-to-Acoustic Model
@@
We also train a semantic-to-acoustic (S2A) model using a masked generative codec transformer conditioned on the semantic tokens. Our semantic-to-acoustic model is based on SoundStorm  (Borsos et al., 2023), which generates multi-layer acoustic token sequences. Given $N$ layers of the acoustic token sequence $\mathbf{A}^{1:N}$, during training, we select one layer $j$ from $1$ to $N$. We denote the $j$th layer of the acoustic token sequence as $A^{j}$. Following the previous discussion, we mask $A^{j}$ at the timestep $t$ to get $\mathbf{A}^{j}_{t}$. The model is then trained to predict $\mathbf{A}^{j}$ conditioned on the prompt $\mathbf{A}^{p}$, the corresponding semantic token sequence $\mathbf{S}$, and all the layers smaller than $j$ of the acoustic tokens. This can be formulated as $p_{\theta_{\text{s2a}}}(\mathbf{A}^{j}|\mathbf{A}^{j}_{t},(\mathbf{A}^{p},\mathbf{S},\mathbf{A}^{1:j-1}))$. We sample $j$ according to a linear schedule $p(j)=1-\frac{2j}{N(N+1)}$. For the input of the S2A model, since the number of frames in the semantic token sequence is equal to the sum of the frames in the prompt acoustic sequence and the target acoustic sequence, we simply sum the embeddings of the semantic tokens and the embeddings of the acoustic tokens from layer $1$ to $j$. During inference, we generate tokens for each layer from coarse to fine, using iterative parallel decoding within each layer. Figure 2 shows a simplified training diagram of the T2S and S2A models.
@@
Figure 2: An overview of training diagram or the T2S (left) and S2A (right) models. The T2S model is trained to predict masked semantic tokens with text and prompt semantic tokens as the prefix. The S2A model is trained to predict masked acoustic tokens of a random layer conditioned on prompt acoustic tokens, semantic tokens, and acoustic tokens of the previous layers.
@@ 3.7 Other Applications
MaskGCT can accomplish tasks beyond zero-shot TTS, such as emotion control, speech content editing, voice conversion, and duration controllable speech translation with simple modifications or the assistance of external tools, demonstrating the potential of MaskGCT as a foundational model for speech generation. We provide more details in Appendix G,  H,  I,  J.
@@ 4.1 Experimental Settings
Datasets. We use the Emilia (He et al., 2024) dataset to train our models. Emilia is a multilingual and diverse in-the-wild speech dataset designed for large-scale speech generation. In this work, we use English and Chinese data from Emilia, each with 50K hours of speech (totaling 100K hours). We evaluate our zero-shot TTS models with three benchmarks: (1) LibriSpeech (Panayotov et al., 2015) test-clean, a widely used test set for English zero-shot TTS. (2) SeedTTS test-en, a test set introduced in Seed-TTS (Anastassiou et al., 2024) of samples extracted from English public corpora, specifically, the set includes 1,000 samples from the Common Voice dataset (Ardila et al., 2019). (3) SeedTTS test-zh, a test set introduced in Seed-TTS of samples extracted from Chinese public corpora, specifically, the set includes 2,000 samples from the DiDiSpeech dataset (Guo et al., 2021).
@@
Evaluation Metrics. We use both objective and subjective metrics to evaluate our models. For the objective metrics, we evaluate speaker similarity (SIM), robustness (WER), and speech quality (FSD). Specifically, for speaker similarity, we compute the cosine similarity between the WavLM TDNN [^1] https://github.com/microsoft/UniSpeech/tree/main/downstreams/speaker_verification (Chen et al., 2022) speaker embedding of generated samples and the prompt. For Word Error Rate (WER), we use a HuBERT-based [^2] https://huggingface.co/facebook/hubert-large-ls960-ft ASR model for LibriSpeech test-clean, Whisper-large-v3 for Seed-TTS test-en, and Paraformer-zh for Seed-TTS test-zh, following previous works. For speech quality, we use Fréchet Speech Distance (FSD) with self-supervised wav2vec 2.0 (Baevski et al., 2020) features, following Le et al. (2024). For the subjective metrics, comparative mean option score (CMOS) and similarity mean option score (SMOS) are used to evaluate naturalness and similarity, respectively. CMOS is on a scale of -3 to 3, and SMOS is on a scale of 1 to 5.
@@
Baseline. We compare our models with state-of-the-art zero-shot TTS systems, including NaturalSpeech 3 (Ju et al., 2024), VALL-E (Wang et al., 2023), VoiceBox (Le et al., 2024), VoiceCraft (Peng et al., 2024), XTTS-v2 (Casanova et al., 2024), and CosyVoice (Du et al., 2024a). More details of each model can be found in Appendix F. We also train an AR T2S model to replace the T2S part of MaskGCT, which can be seen as the SoundStorm model for TTS.
@@
Training. We train all models on 8 NVIDIA A100 80GB GPUs. We train two T2S models of different sizes (denoted as T2S-Base and T2S-large). For more details about the model architecture, please refer to Appendix A. We report the metrics of T2S-large by default, and you can find a comparison of model sizes in Section 4.3. We also compare two different methods of text tokenization: Grapheme-to-Phoneme(G2P) (Bernard and Titeux, 2021) and Byte Pair Encoding (BPE) (Gage, 1994). See more details of the two methods in Appendix D. We report the metrics of G2P by default. We optimize these models with the AdamW (Loshchilov and Hutter, 2017) optimizer with a learning rate of 1e-4 and 32K warmup steps, following the inverse square root learning schedule. We use the classifier-free guidance (Ho and Salimans, 2022), during training for both the T2S and S2A models, we drop the prompt with a probability of 0.15. See more details about classifier-free guidance and classifier-free guidance rescale in Appendix E.
@@
Inference. For the T2S model, we use 50 steps as the default total inference steps. The classifier-free guidance scale and the classifier-free guidance rescale factor (Lin et al., 2024) are set to 2.5 and 0.75, respectively. For sampling, we use a top-k of 20, with the sampling temperature annealing from 1.5 to 0. We add Gumbel noise to token confidences when determining the remasking process, following Chang et al. (2022). For the S2A model, we use $[40,16,1,1,1,1,1,1,1,1,1,1]$ steps for acoustic RVQ layers by default, we find the S2A model can also perform well with fewer inference steps of $[10,1,1,1,1,1,1,1,1,1,1,1]$. We use the same sampling strategy as the T2S model, except that we use greedy sampling instead of top-k sampling if the inference step is 1.
@@ 4.2 Zero-Shot TTS
Ground Truth 0.68 1.94 - 4.048 0.000
VALL-E (Wang et al., 2023) 0.50 5.90 - 3.472 -0.522
VoiceBox (Le et al., 2024) 0.64 2.03 0.762 3.802 -0.407
NaturalSpeech 3 (Ju et al., 2024) 0.67 1.94 0.786 4.258 0.164
VoiceCraft (Peng et al., 2024) 0.45 4.68 0.981 3.519 -0.333
XTTS-v2 (Casanova et al., 2024) 0.51 4.20 0.945 3.015 -0.985
MaskGCT 0.687(0.723) 2.634(1.976) 0.886 4.274 0.109
MaskGCT (gt length) 0.697 2.012 0.746 4.333 0.130
@@
Ground Truth 0.730 2.143 - 3.923 0.000
CosyVoice (Du et al., 2024a) 0.643 4.079 0.316 3.516 -0.416
XTTS-v2 (Casanova et al., 2024) 0.463 3.248 0.484 3.150 -0.866
VoiceCraft (Peng et al., 2024) 0.470 7.556 0.226 3.179 -1.083
MaskGCT 0.717(0.760) 2.623(1.283) 0.188 4.242 0.033
MaskGCT (gt length) 0.728 2.466 0.159 4.129 0.122
@@
Ground Truth 0.750 1.254 - 3.863 0.000
CosyVoice (Du et al., 2024a) 0.750 4.089 0.276 3.544 -0.450
XTTS-v2 (Casanova et al., 2024) 0.635 2.876 0.413 2.950 -0.807
MaskGCT 0.774(0.805) 2.273(0.843) 0.106 4.094 0.050
MaskGCT (gt length) 0.777 2.183 0.101 4.107 0.083
In this section, we show the main results of zero-shot TTS: we show comparison results with SOTA baselines in Section 4.2.1; we compare MaskGCT with replacing T2S model to an AR model in Section 4.2.2; We present the performance of MaskGCT across varying speech tempos in Section 4.2.3.

### v2

@@ 3.2.1 Speech Semantic Representation Codec
Discrete speech representations can be divided into semantic tokens and acoustic tokens. Generally, semantic tokens are obtained by discretizing features from speech self-supervised learning (SSL). Previous two-stage, large-scale TTS systems Betker (2023); Borsos et al. (2023a); Kharitonov et al. (2023) typically first use text to predict semantic tokens, and then employ another model to predict acoustic tokens or features. This is because semantic tokens have a stronger correlation with text or phonemes, which makes predicting them more straightforward than directly predicting acoustic tokens. Commonly, previous works have used k-means to discretize semantic features to obtain semantic tokens; however, this method can lead to a loss of information. This loss may complicate the accurate reconstruction of high-quality speech or the precise prediction of acoustic tokens, especially for tonally rich languages. For example, our early experiments demonstrate the challenges of accurately predicting acoustic tokens to achieve proper prosody for Chinese using semantic tokens obtained via k-means. Therefore, we need to discretize semantic representation features while minimizing information loss. Inspired by Huang et al. (2023), we train a VQ-VAE model to learn a vector quantization codebook that reconstructs speech semantic representations from a speech SSL model. For a speech semantic representation sequence $\mathbf{S}\in\mathbb{R}^{T\times d}$, the vector quantizer quantizes the output of the encoder $\mathcal{E}(\mathbf{S})$ to $\mathbf{E}$, and the decoder reconstructs $\mathbf{E}$ back to $\hat{\mathbf{S}}$. We optimize the encoder and the decoder using a reconstruction loss between $\mathbf{S}$ and $\hat{\mathbf{S}}$, employ codebook loss to optimize the codebook and use commitment loss to optimize the encoder with the straight-through method Van Den Oord et al. (2017). The total loss for training the semantic representation codec can be written as:
@@
In detail, we utilize the hidden states from the 17th layer of W2v-BERT 2.0 Chung et al. (2021) as the semantic features for our speech encoder. The encoder and decoder are composed of multiple ConvNext Liu et al. (2022a) blocks. Following the methods of improved VQ-GAN Yu et al. (2021) and DAC Kumar et al. (2024), we use factorized codes to project the output of the encoder into a low-dimensional latent variable space. The codebook contains 8,192 entries, each of dimension 8. Further details about the model architecture are provided in Appendix A.4.
3.2.2 Text-to-Semantic Model
@@
Based on the previous discussion, we employ a non-autoregressive masked generative transformer to train a text-to-semantic (T2S) model, instead of using an autoregressive model or any text-to-speech alignment information. During training, we randomly extract a portion of the prefix of the semantic token sequence as the prompt, denoted as $\mathbf{S}^{p}$. We then concatenate the text token sequence $\mathbf{P}$ with $\mathbf{S}^{p}$ to form the condition. We simply add $(\mathbf{P},\mathbf{S}^{p})$ as the prefix sequence to the input masked semantic token sequence $\mathbf{S}_{t}$ to leverage the in-context learning ability of language models. We use a Llama-style Touvron et al. (2023) transformer as the backbone of our model, incorporating gated linear units with GELU Hendrycks and Gimpel (2016) activation, rotation position encoding Su et al. (2024), etc., but replacing causal attention with bidirectional attention. We also use adaptive RMSNorm Zhang and Sennrich (2019), which accepts the time step $t$ as the condition.
@@
During inference, we generate the target semantic token sequence of any specified length conditioned on the text and the prompt semantic token sequence. In this paper, we also train a flow matching Lipman et al. (2022) based duration prediction model to predict the total duration conditioned on the text and prompt speech duration, leveraging in-context learning. More details can be found in Appendix A.5.
3.2.3 Semantic-to-Acoustic Model
@@
We also train a semantic-to-acoustic (S2A) model using a masked generative codec transformer conditioned on the semantic tokens. Our semantic-to-acoustic model is based on SoundStorm  Borsos et al. (2023a), which generates multi-layer acoustic token sequences. Given $N$ layers of the acoustic token sequence $\mathbf{A}^{1:N}$, during training, we select one layer $j$ from $1$ to $N$. We denote the $j$th layer of the acoustic token sequence as $A^{j}$. Following the previous discussion, we mask $A^{j}$ at the timestep $t$ to get $\mathbf{A}^{j}_{t}$. The model is then trained to predict $\mathbf{A}^{j}$ conditioned on the prompt $\mathbf{A}^{p}$, the corresponding semantic token sequence $\mathbf{S}$, and all the layers smaller than $j$ of the acoustic tokens. This can be formulated as $p_{\theta_{\text{s2a}}}(\mathbf{A}^{j}|\mathbf{A}^{j}_{t},(\mathbf{A}^{p},\mathbf{S},\mathbf{A}^{1:j-1}))$. We sample $j$ according to a linear schedule $p(j)=1-\frac{2j}{N(N+1)}$. For the input of the S2A model, since the number of frames in the semantic token sequence is equal to the sum of the frames in the prompt acoustic sequence and the target acoustic sequence, we simply sum the embeddings of the semantic tokens and the embeddings of the acoustic tokens from layer $1$ to $j$. During inference, we generate tokens for each layer from coarse to fine, using iterative parallel decoding within each layer. Figure 2 shows a simplified training diagram of the T2S and S2A models.
@@
Figure 2: An overview of training diagram of the T2S (left) and S2A (right) models. The T2S model is trained to predict masked semantic tokens with text and prompt semantic tokens as the prefix. The S2A model is trained to predict masked acoustic tokens of a random layer conditioned on prompt acoustic tokens, semantic tokens, and acoustic tokens of the previous layers.
3.2.4 Speech Acoustic Codec
Speech acoustic codec is trained to quantize speech waveform to multi-layer discrete tokens while aiming to preserve all the information of the speech as soon as possible. We follow the residual vector quantization (RVQ) method to compress the 24K sampling rate speech waveform into discrete tokens of 12 layers. The codebook size of each layer is 1,024 and the codebook dimension is 8. The model architectures, discriminators, and training losses follow DAC Kumar et al. (2024), except that we use the Vocos Siuzdak (2023) architecture as the decoder for more efficient training and inference. Figure 5 shows the comparison between the semantic codec and acoustic codec.
@@ 3.3 Other Applications
MaskGCT can accomplish tasks beyond zero-shot TTS, such as duration-controllable speech translation (cross-lingual dubbing), emotion control, speech content editing, and voice conversion with simple modifications or the assistance of external tools, demonstrating the potential of MaskGCT as a foundational model for speech generation. We provide more details in Appendix F,  G,  H,  I.
@@ 4.1 Experimental Settings
Datasets. We use the Emilia He et al. (2024) dataset to train our models. Emilia is a multilingual and diverse in-the-wild speech dataset designed for large-scale speech generation. In this work, we use English and Chinese data from Emilia, each with 50K hours of speech (totaling 100K hours). We evaluate our zero-shot TTS models with three benchmarks: (1) LibriSpeech Panayotov et al. (2015) test-clean, a widely used test set for English zero-shot TTS. (2) SeedTTS test-en, a test set introduced in Seed-TTS Anastassiou et al. (2024) of samples extracted from English public corpora, includes 1,000 samples from the Common Voice dataset Ardila et al. (2019). (3) SeedTTS test-zh, a test set introduced in Seed-TTS of samples extracted from Chinese public corpora, includes 2,000 samples from the DiDiSpeech dataset Guo et al. (2021). We also scale the training dataset to six languages to support multilingual zero-shot TTS. We provide additional experimental details and evaluation results about multilingual zero-shot TTS in Appendix E.
@@
Evaluation Metrics. We use both objective and subjective metrics to evaluate our models. For the objective metrics, we evaluate speaker similarity (SIM-O), robustness (WER), and speech quality (FSD). Specifically, for speaker similarity, we compute the cosine similarity between the WavLM TDNN [^2] https://github.com/microsoft/UniSpeech/tree/main/downstreams/speaker_verification Chen et al. (2022) speaker embedding of generated samples and the prompt. For Word Error Rate (WER), we use a HuBERT-based [^3] https://huggingface.co/facebook/hubert-large-ls960-ft ASR model for LibriSpeech test-clean, Whisper-large-v3 for Seed-TTS test-en, and Paraformer-zh for Seed-TTS test-zh, following previous works. For speech quality, we use Fréchet Speech Distance (FSD) with self-supervised wav2vec 2.0 Baevski et al. (2020) features, following Le et al. (2024). For the subjective metrics, comparative mean option score (CMOS) and similarity mean option score (SMOS) are used to evaluate naturalness and similarity, respectively. CMOS is on a scale of -3 to 3, and SMOS is on a scale of 1 to 5.
@@
Baseline. We compare our models with state-of-the-art zero-shot TTS systems, including NaturalSpeech 3 Ju et al. (2024), VALL-E Wang et al. (2023), VoiceBox Le et al. (2024), VoiceCraft Peng et al. (2024), XTTS-v2 Casanova et al. (2024), and CosyVoice Du et al. (2024a). More details of each model can be found in Appendix D. We also train an AR-based T2S model to replace the T2S part of MaskGCT, we term it as AR + SoundStorm.
@@
Training. We train all models on 8 NVIDIA A100 80GB GPUs. We train two T2S models of different sizes (denoted as T2S-Base and T2S-large). For more details about the model architecture, please refer to Appendix A.1. We report the metrics of T2S-large by default, and you can find a comparison of model sizes in Section 4.4. We also compare two different methods of text tokenization: Grapheme-to-Phoneme (G2P) Bernard and Titeux (2021) and Byte Pair Encoding (BPE) Gage (1994). See more details of the two methods in Appendix A.6. We report the metrics of G2P by default. We optimize these models with the AdamW Loshchilov and Hutter (2017) optimizer with a learning rate of 1e-4 and 32K warmup steps, following the inverse square root learning schedule. We use the classifier-free guidance Ho and Salimans (2022), during training for both the T2S and S2A models, we drop the prompt with a probability of 0.15. See more details about classifier-free guidance and classifier-free guidance rescale in Appendix C.
@@
Inference. For the T2S model, we use 50 steps as the default total inference steps. The classifier-free guidance scale and the classifier-free guidance rescale factor Lin et al. (2024) are set to 2.5 and 0.75, respectively. For sampling, we use a top-k of 20, with the sampling temperature annealing from 1.5 to 0. We add Gumbel noise to token confidences when determining the remasking process, following Chang et al. (2022). For the S2A model, we use $[40,16,1,1,1,1,1,1,1,1,1,1]$ steps for acoustic RVQ layers by default, we find the S2A model can also perform well with fewer inference steps of $[10,1,1,1,1,1,1,1,1,1,1,1]$ (see Appendix A.3). We use the same sampling strategy as the T2S model, except that we use greedy sampling instead of top-k sampling if the inference step is 1.
@@ 4.2 Zero-Shot TTS
Ground Truth 0.68 1.94 - 4.05±${}_{\scriptscriptstyle\pm\text{0.12}}$ 0.00
VALL-E Wang et al. (2023) 0.50 5.90 - 3.47±${}_{\scriptscriptstyle\pm\text{0.26}}$0.52±${}_{\scriptscriptstyle\pm\text{0.22}}$
VoiceBox Le et al. (2024) 0.64 2.03 0.762 3.80±${}_{\scriptscriptstyle\pm\text{0.17}}$0.41±${}_{\scriptscriptstyle\pm\text{0.13}}$
NaturalSpeech 3 Ju et al. (2024) 0.67 1.94 0.786 4.26±${}_{\scriptscriptstyle\pm\text{0.10}}$ 0.16±${}_{\scriptscriptstyle\pm\text{0.14}}$
VoiceCraft Peng et al. (2024) 0.45 4.68 0.981 3.52±${}_{\scriptscriptstyle\pm\text{0.21}}$0.33±${}_{\scriptscriptstyle\pm\text{0.16}}$
XTTS-v2 Casanova et al. (2024) 0.51 4.20 0.945 3.02±${}_{\scriptscriptstyle\pm\text{0.22}}$0.98±${}_{\scriptscriptstyle\pm\text{0.19}}$
MaskGCT 0.687(0.723) 2.634(1.976) 0.886 4.27±${}_{\scriptscriptstyle\pm\text{0.14}}$ 0.10±${}_{\scriptscriptstyle\pm\text{0.16}}$
MaskGCT (gt length) 0.697 2.012 0.746 4.33±${}_{\scriptscriptstyle\pm\text{0.11}}$ ${}_{\scriptscriptstyle\pm\text{0.13}}$
@@
Ground Truth 0.730 2.143 - 3.92±${}_{\scriptscriptstyle\pm\text{0.15}}$ 0.00
CosyVoice Du et al. (2024a) 0.643 4.079 0.316 3.52±${}_{\scriptscriptstyle\pm\text{0.17}}$0.41±${}_{\scriptscriptstyle\pm\text{0.18}}$
XTTS-v2 Casanova et al. (2024) 0.463 3.248 0.484 3.15±${}_{\scriptscriptstyle\pm\text{0.22}}$0.86±${}_{\scriptscriptstyle\pm\text{0.19}}$
VoiceCraft Peng et al. (2024) 0.470 7.556 0.226 3.18±${}_{\scriptscriptstyle\pm\text{0.20}}$1.08±${}_{\scriptscriptstyle\pm\text{0.15}}$
MaskGCT 0.717(0.760) 2.623(1.283) 0.188 4.24±${}_{\scriptscriptstyle\pm\text{0.12}}$ 0.03±${}_{\scriptscriptstyle\pm\text{0.14}}$
MaskGCT (gt length) 0.728 2.466 0.159 4.13±${}_{\scriptscriptstyle\pm\text{0.17}}$ 0.12±${}_{\scriptscriptstyle\pm\text{0.15}}$
@@
Ground Truth 0.750 1.254 - 3.86±${}_{\scriptscriptstyle\pm\text{0.17}}$ 0.00
CosyVoice Du et al. (2024a) 0.750 4.089 0.276 3.54±${}_{\scriptscriptstyle\pm\text{0.12}}$0.45±${}_{\scriptscriptstyle\pm\text{0.15}}$
XTTS-v2 Casanova et al. (2024) 0.635 2.876 0.413 2.95±${}_{\scriptscriptstyle\pm\text{0.18}}$0.81±${}_{\scriptscriptstyle\pm\text{0.22}}$
MaskGCT 0.774(0.805) 2.273(0.843) 0.106 4.09±${}_{\scriptscriptstyle\pm\text{0.12}}$ 0.05±${}_{\scriptscriptstyle\pm\text{0.17}}$
MaskGCT (gt length) 0.777 2.183 0.101 4.11±${}_{\scriptscriptstyle\pm\text{0.12}}$ 0.08±${}_{\scriptscriptstyle\pm\text{0.18}}$
In this section, we show the main results of zero-shot TTS: we show comparison results with SOTA baselines in Section 4.2.1; we compare MaskGCT with replacing T2S model to an AR model in Section 4.2.2; We present the performance of MaskGCT across varying speech tempos in Section 4.2.3. Additionally, we present the results of zero-shot TTS for speech style imitation in Section 4.3, multilingual zero-shot TTS in Appendix E, and cross-lingual speech translation (dubbing) in Appendix F.